OpenStudy (starryknight):

Find the sum of the first 26 terms of the following sequence: 12, 18, 24 . . .

OpenStudy (starryknight):

i got 2210 but idont know if its right or wrong

OpenStudy (kirbykirby):

Hmm I get 2262

OpenStudy (kirbykirby):

\[\large \sum_{n=0}^{25}(12+6n)=\sum_{n=1}^{26}(6+6n)=6(26)+6\frac{26(26+1)}{2} \]